# Avoiding matrix exponentials for large transition rate matrices

### 09 Dec 2023 v1

@@ Abstract
Exact methods for exponentiation of matrices of dimension $N$ can be computationally expensive in terms of execution time ($N^{3}$) and memory requirements ($N^{2}$) not to mention numerical precision issues. A type of matrix often exponentiated in the sciences is the rate matrix. Here we explore five methods to exponentiate rate matrices some of which apply even more broadly to other matrix types. Three of the methods leverage a mathematical analogy between computing matrix elements of a matrix exponential and computing transition probabilities of a dynamical processes (technically a Markov jump process, MJP, typically simulated using Gillespie). In doing so, we identify a novel MJP-based method relying on restricting the number of “trajectory” jumps based on the magnitude of the matrix elements with favorable computational scaling. We then discuss this method’s downstream implications on mixing properties of Monte Carlo posterior samplers. We also benchmark two other methods of matrix exponentiation valid for any matrix (beyond rate matrices and, more generally, positive definite matrices) related to solving differential equations: Runge-Kutta integrators and Krylov subspace methods. Under conditions where both the largest matrix element and the number of non-vanishing elements scale linearly with $N$ — reasonable conditions for rate matrices often exponentiated — computational time scaling with the most competitive methods (Krylov and one of the MJP-based methods) reduces to $N^{2}$ with total memory requirements of $N$.
@@ 1 Introduction
where $\bar{\rho}(t)$ is a (row) vector of probabilities for the system’s states and $\mathbf{A}_{\theta}$ is the generator matrix also called a transition rate matrix (where all rows sum to one) whose elements are kinetic parameters. We collect the unique set of kinetic parameters that make up $\mathbf{A}_{\theta}$ in $\theta$.
@@
In fact, in modeling inspired from chemical systems, whose acceptable state spaces vastly exceed the number of reactions allowed, we often encounter sparse generator matrices. Indeed, properly exploiting sparse matrices provides an immediate cost benefit since storing them requires memory scaling with the number of non-zero elements, i.e., linearly with $N$ or better. Furthermore, avoiding dense matrix exponentiation using only sparse matrix-vector products incurs an additional benefit to computational time scaling, reducing the total computational cost to scale with the number of non-zero elements, again, often linearly or better with $N$.
@@
The first MJP-based method (#1), which we refer to as trajectory MJP-based method (T-MJP) was developed by Rao and Teh in Ref. [28]. In T-MJP the system’s trajectory is sampled through the forward-filtering backward-sampling algorithm [28, 31, 15, 3, 29]. This method learns the parameters $\theta$ in a manner that is statistically equivalent to exactly solving (1). As we will show, T-MJP’s trajectory sampling (as a substitute for matrix exponentiation) time scaling is $N^{2}$ (an improvement over $N^{3}$ for naive matrix exponential), but unfortunately retains the $N^{2}$ memory scaling described in (ii). Even worse, since T-MJP avoids the matrix exponential by introducing (de-marginalizing over) a large number of latent random variables (i.e., the trajectories which includes the jump times and states visited after each jump), it severely increases the dimensionality of any statistical inference task. As a result, many MCMC iterations are needed to sample both trajectories and parameters. That is, the MCMC chains suffer from “poor mixing” giving rise to slow convergence. These computational issues may be avoided if trajectory samples are not needed.
@@
Method #3, referred to as remarginalized MJP-based method (R-MJP), improves mixing by entirely marginalizing over trajectories. Although marginalizing over trajectories demands an approximation not required by Methods #1 and #2, R-MJP provides an upper bound in the total vector times matrix exponentiation error, with no additional computation [30]. With the same scaling as both previous methods, we will demonstrate R-MJP’s considerably better mixing than Methods #1 and #2.
The two remaining methods treat the solution of (1) without leveraging (i.e., de-marginalizing over) trajectories, or indeed any facet of MJP structure. We therefore refer to them as direct probability calculation methods.
Method #4 simply applies the Runge-Kutta method for solving differential equations [32] to avoid matrix exponentiation. Method #4 proceeds by slicing the time projection into small intervals of size $\Delta t$ and evolving (1) within that time interval. Naturally the choice of $\Delta t$ is arbitrary, thus making it significantly harder to keep track of the upper bound in the approximation error. Thus $\bar{\rho}(t)$ arises from the sequence $,\{\bar{\rho(0)},\bar{\rho(\Delta t)},\bar{\rho(2\Delta t)},$$\ldots$,… , $\bar{\rho(t)}\}$. We focus in particular on the fourth-order Runge-Kutta methods (RK4), an implementation typically chosen as a balance between precision and computational time[32]. As usual with RK4, time scaling is $N^{2}$ and memory requirements scale as $N$.
@@
In what follows, we implement methods #1-#5 comparing their computational time scaling, memory requirements, and mixing. In short, we want to observe which method is best to create a MCMC chain that runs sufficient iterations in reasonable time. Due to method #1’s excessive memory requirements, we avoid comparing it to other methods and proceed by benchmarking the remaining methods (#2-#5). We show that when compared to RK4 and R-MJP, J-MJP requires a much larger number of matrix exponential equivalent calculations to complete a modeling task, as predicted. From our benchmark, we conclude that either R-MJP (Method #3) or Krylov subspace (Method #5) attains the fastest overall computational time, depending on both the state space and the dynamics at hand. A diagram illustrating the methods and the reasons some are considered unsuitable are shown in Fig. 1 c).
@@ 2 Methods
2.1 Necessary background on MJPs
Here we describe a general theory of continuous time systems transitioning between distinct states, $\sigma_{n}$ [1, 2, 3, 29]. We collect any system’s states into a set termed the state space $S=\{\sigma_{0},\sigma_{1},\ldots,\sigma_{N-1}\}$. We describe the transition dynamics within the state space by a transition rate matrix $\mathbf{\Lambda}$ whose elements $\lambda_{nm}$, called transition rates from state $\sigma_{n}$ to $\sigma_{m}$, imply that the probability of the system transitioning from $\sigma_{n}$ to $\sigma_{m}$ in an infinitesimal time interval $\mathrm{d}t$ is $\lambda_{nm}\mathrm{d}t$ [3]. By definition, all self-transition rates are zero, $\lambda_{nn}=0\ \forall n$, leaving $N(N-1)$ potentially independent parameters.
In many problems of scientific interest, some examples are given in Sec. 3.1, we will see that the remaining $\lambda_{nm}$ are typically described using a much smaller number of unique parameters, collected under $\theta$, from which $\lambda_{nm}$ are reconstructed. For example, in a simple death process, the rates in going from $n$ members to $n-1$ members, $\lambda_{n,n-1}$, are all the same irrespective of $n$ (and coincide with the death rate). To make $\theta$ dependence explicit for matrices, we define $\mathbf{\Lambda}_{\theta}=\mathbf{\Lambda}(\theta)$ and denote the count of non-zero elements in a matrix $\mathbf{M}$ as $[\mathbf{M}]$. For example, for a simple death process starting from $N$ members, we have $N$ non-zero elements.
Such systems in discrete space evolving in continuous time with rate matrices are termed MJPs. An MJP’s dynamics in a time interval $[0,T]$, usually thought of as an experimental time-course, are described completely by the states the system occupies over time, $\tilde{s}_{0},\tilde{s}_{1},\ldots,\tilde{s}_{l}$ with each $\tilde{s}_{l^{\prime}}\in S$, and the time at which each transition occurs, $t_{1},t_{2},\ldots,t_{l}$. Together, these describe a trajectory, an equivalent description of the MJP’s dynamics as function $\mathcal{T}(\cdot)$ mapping times from the interval $[0,T]$ onto the state space $S$
@@ 2.1.1 Forward Modeling
Now, sampling trajectories from (10) yields MJP trajectories mediated by $\mathbf{A}_{\theta}$, described in Sec 2.1. As proven in Ref. [27], the jump times themselves may be uniformly sampled across $[0,t]$, resulting in using the term “uniformization” to describe the use of (10) in the simulation of MJPs [27, 28, 3]. Unlike the Gillespie simulation, $k$ must include self-transitions for uniformization to remain equivalent to its corresponding MJP under uniform jump time statistics. We therefore follow the uniformization literature and say $k$ counts ‘virtual jumps’ (i.e. it is distinct from the number of transitions $l$ in the previous section), to indicate that it includes both real and self-transitions. Although uniformization is unpopular relative to Gillespie’s more efficient stochastic simulation algorithm [35], we describe it here since it allows us to propose trajectories conditioned on data and subsequently accept or reject them within Monte Carlo.
@@ 2.1.2 Inverse Modeling
With the details of forward modeling established, we turn to the task, the so-called inverse problem, of obtaining the probability of the state $\rho_{n}$, from $\theta$, usually requiring matrix exponentiation. In the inverse problem, information on $\theta$ is provided through observations of MJP states. Notably, Methods #1 and #2 provide stochastic samples of trajectories and jumps, respectively, contributing to the matrix exponential while Methods #3 – #5 , in principle, could be used to directly compute the vector times matrix exponential directly as in (2). Later, Sec. 3.2, we will embed all the methods within an inverse modeling scheme as a means of assessing the efficiency of each method.
Now we consider multiple back to back intervals, $[0,T]$. We collect the end times of each interval as the set $\bar{T}=\{T^{1},T^{2},\ldots,T^{I}\}$. The state of a system could be observed at these times leading to a sequence of state observations $\bar{s}=\{s^{1},s^{2},\ldots,s^{I}\}$. Alternately, we could also consider a different system drawn from a collection of systems at times $\bar{T}=\{T^{1},T^{2},\ldots,T^{I}\}$ and measure its state at those times also yielding a sequence of state observations $\bar{s}=\{s^{1},s^{2},\ldots,s^{I}\}$. The latter falls within the realm of what is sometimes called snapshot data. The methods we discuss here apply to both type of observations though, for ease of computation, our benchmarking of Methods #1 – #5 is performed on snapshot data.
Following the notation established in (3), from the state observations we can write we write $p(s^{i}|\theta)=p(\mathcal{T}^{i}(T^{i})=s^{i}|\theta)$, often termed the likelihood. As defined in (4), we have $p(s^{i}|\theta)=\rho(T^{i})_{s^{i}}=(\bar{\rho}(0)e^{\mathbf{A}_{\theta}T^{i}})_{s^{i}}$ thus equivalent to what is obtained by the vector times matrix exponential in (2).
In the following section we will explain the different ways to avoid the full matrix exponential calculation, but first, we establish the notation necessary to explain Methods #1 – #3. Learning the kinetic parameters $\theta$ from the data means obtaining $p(\theta|\bar{s})$ which is related to likelihood through Bayes’ theorem
@@
For Methods #1 and #2, we avoid calculating $p(\theta|\bar{s})$ by making such calculations based on latent variables $\mathcal{T}^{i}$, the trajectories, and $k^{i}$, the number of virtual jumps, respectively. When these latent variables from are used for sampling, we simply ‘demarginalize’ them from (11). That is, we re-write the posterior as the following from marginalization over trajectories and virtual jumps, respectively,
@@ 2.2 Avoiding the matrix exponential in Methods #1 – #5
We obtain compact scaling relationships between different methods under two assumptions inspired by relevant physical examples: a) the number of nonzero elements in $\mathbf{A}_{\theta}$ scales linearly with $N$ (conveniently guaranteeing sparse $\mathbf{A}_{\theta}$) and b) the exponent of any power of $N$ appearing explicitly in $\mathbf{A}_{\theta}$ is less than or equal to $1$ as larger rates can introduce more jumps in the trajectories introduced in Methods #1 – #3 and reduce the time step of integration of Methods #4 and #5. To be clear, both conditions are used to compute scaling laws, but not required for the methods to work.
Regardless of which method we are discussing, when dealing with chemical reactions, the number of nonzero elements scales at worst linearly with the state space size, $O([\mathbf{A}_{\theta}])=O(N)$, as in condition a). It is also usually the case that when molecular species do not interact with themselves, the largest element of the rate matrix scales linearly with the state space — $O(\max\limits_{n}{|a_{nn}|})=O(N)$, as in condition b). As we will see, all methods have some dependence on the largest value in the rate matrix, $\max\limits_{n}{|a_{nn}|}$. As this quantity represents the fastest time-scale at which dynamics occur, consequently, all methods must adapt to it in order to provide accurate results.
Table 1: Table summarizing the methods discussed, their number of iterations (first column) and time cost per iteration (second column), with $\kappa$ being the number of vectors taken into the Krylov subspace. The third column (time scaling) is built by multiplying the first and second columns, assuming conditions a) and b) outlined in the second paragraph of Sec. 2.2. Although J-MJP appears competitive with other methods, we will see in Sec. 3.2.2 that the latent variables it requires lead to poor MCMC mixing.

### 06 Feb 2024 v3

@@ Abstract
Exact methods for exponentiation of matrices of dimension $N$ can be computationally expensive in terms of execution time ($N^{3}$) and memory requirements ($N^{2}$) not to mention numerical precision issues. A matrix often exponentiated in the natural sciences is the rate matrix. Here we explore five methods to exponentiate rate matrices some of which apply more broadly to other matrix types. Three of the methods leverage a mathematical analogy between computing matrix elements of a matrix exponential and computing transition probabilities of a dynamical processes (technically a Markov jump process, MJP, typically simulated using Gillespie). In doing so, we identify a novel MJP-based method relying on restricting the number of “trajectory” jumps that incurs improved computational scaling. We then discuss this method’s downstream implications on mixing properties of Monte Carlo posterior samplers. We also benchmark two other methods of matrix exponentiation valid for any matrix (beyond rate matrices and, more generally, positive definite matrices) related to solving differential equations: Runge-Kutta integrators and Krylov subspace methods. Under conditions where both the largest matrix element and the number of non-vanishing elements scale linearly with $N$ — reasonable conditions for rate matrices often exponentiated — computational time scaling with the most competitive methods (Krylov and one of the MJP-based methods) reduces to $N^{2}$ with total memory requirements of $N$.
@@ 1 Introduction
where $\bar{\rho}(t)$ is a (row) vector of probabilities for the system’s states and $\mathbf{A}_{\theta}$ is the generator matrix also called a transition rate matrix (where all rows sum to one). Usually, the elements of $\mathbf{A}_{\theta}$ are functions of small number of kinetic parameters – e.g., the propensity of each chemical reaction modelled. We denote this concise set of unique parameters as $\theta$. To illustrate this with practical applications, Sec. 3.1 provides specific examples.
@@
In fact, in modeling inspired from chemical systems, whose acceptable state spaces vastly exceed the number of reactions allowed, we often encounter sparse generator matrices. Indeed, properly exploiting sparse matrices provides an immediate cost benefit since storing them requires memory scaling with the number of non-zero elements, i.e., linear scaling or better with $N$. Furthermore, since for irreducible dynamics the matrix exponential will be fully dense, avoiding dense matrix exponentiation using only sparse matrix-vector products incurs an additional benefit to computational time scaling, reducing the total computational cost to scale with the number of non-zero elements, which even beyond modeling chemical systems, often scale linearly or better with $N$.
@@
The first MJP-based method (#1), which we refer to as trajectory MJP-based method (T-MJP) was developed by Rao and Teh in Ref. [28]. In T-MJP the system’s trajectory is sampled through the forward-filtering backward-sampling algorithm [28, 31, 15, 3, 29]. This method learns the parameters $\theta$ in a manner that is statistically equivalent to exactly solving (1). As we will show, T-MJP’s trajectory sampling (as a substitute for matrix exponentiation) time scaling is $N^{2}$ (an improvement over $N^{3}$ for naive matrix exponential), but unfortunately retains the $N^{2}$ memory scaling described in (ii). Even worse, since T-MJP avoids the matrix exponential by introducing a large number of latent random variables (i.e., the trajectories which includes the jump times and states visited after each jump), it severely increases the dimensionality of any statistical inference task. As a result, many MCMC iterations are needed to sample both trajectories and parameters. That is, the MCMC chains suffer from “poor mixing” giving rise to slow convergence. These computational issues may be avoided if trajectory samples are not needed.
@@
Method #3, referred to as remarginalized MJP-based method (R-MJP), improves mixing by entirely marginalizing over trajectories, meaning integrating over all trajectories ending in the same observation. Although marginalizing over trajectories demands an approximation not required by Methods #1 and #2, R-MJP provides an upper bound in the total error in the final result, with no additional computation [30]. With the same scaling as both previous methods, we will demonstrate R-MJP’s considerably better mixing than Methods #1 and #2.
The two remaining methods treat the solution of (1) without leveraging trajectories, or indeed any facet of MJP structure. We therefore refer to them as direct probability calculation methods.
Method #4 simply applies the Runge-Kutta method for solving differential equations [32] to avoid matrix exponentiation. Method #4 proceeds by slicing the time projection into small intervals of size $\Delta t$ and evolving (1) within that time interval. Naturally the choice of $\Delta t$ is arbitrary, making it significantly harder to keep track of the upper bound in the approximation error. Thus $\bar{\rho}(t)$ arises from the sequence $,\{\bar{\rho(0)},\bar{\rho(\Delta t)},\bar{\rho(2\Delta t)},$$\ldots$,… , $\bar{\rho(t)}\}$. We focus in particular on the fourth-order Runge-Kutta methods (RK4), an implementation typically chosen as a balance between precision and computational time[32]. As usual with RK4, time scaling is $N^{2}$ and memory requirements scale as $N$.
@@
In what follows, we implement methods #1-#5 comparing their computational time scaling, memory requirements, and mixing. We do so in order to determine which method is best suited for iterative MCMC aimed at estimating parameters and their associated uncertainty. Due to method #1’s excessive memory requirements, we avoid comparing it to other methods and proceed by benchmarking the remaining methods (#2-#5). We show that when compared to RK4 and R-MJP, J-MJP requires a much larger number of matrix exponential equivalent calculations to complete a modeling task, as predicted. From our benchmark, we conclude that either R-MJP (Method #3) or Krylov subspace (Method #5) attains the fastest overall computational time, depending on both the state space and the dynamics at hand. A guide for when to use each method is given in the conclusion (Sec. 4). A diagram illustrating the methods and the reasons some are considered unsuitable are shown in Fig. 1 c).
@@ 2 Methods
This section outlines two key points: first, as an overview, we introduce the necessary background on Markovian dynamics within discrete state spaces in Sec. 2.1. Next, in Sec. 2.2, we elaborate on the key aspects required in implementing Methods # 1 – # 5. While technical details presented herein are essential for those intending to immediately apply these methods, readers primarily interested in application and benchmarking may directly proceed to Sec. 3. Moreover, code for the implementation of methods #2 – #5 (using the Python library for compiled sparse matrices [35]) can be found in our GitHub repository [36].
2.1 Background on MJPs
Here we describe a general theory of continuous time systems transitioning between distinct states, $\sigma_{n}$ [1, 2, 3, 29]. We collect any system’s states into a set termed the state space $\sigma_{1:N}=\{\sigma_{0},\sigma_{1},\ldots,\sigma_{N-1}\}$ We describe the transition dynamics within the state space by a transition rate matrix $\mathbf{\Lambda}$ whose elements $\lambda_{nm}$, called transition rates from state $\sigma_{n}$ to $\sigma_{m}$, imply that the probability of the system transitioning from $\sigma_{n}$ to $\sigma_{m}$ in an infinitesimal time interval $\mathrm{d}t$ is $\lambda_{nm}\mathrm{d}t$. By definition, all self-transition rates are zero, $\lambda_{nn}=0\ \forall n$, leaving $N(N-1)$ potentially independent parameters.
In many problems of scientific interest, of which some examples are provided in Sec. 3.1, we will see that the remaining $\lambda_{nm}$ are typically described using a much smaller number of unique parameters, collected under $\theta$, from which $\lambda_{nm}$ are determined. For example, in a simple death process, the rates in going from $n$ members to $n-1$ members, $\lambda_{n,n-1}$, are all the same irrespective of $n$ (and coincide with the death rate). To make $\theta$ dependence explicit for matrices, we define $\mathbf{\Lambda}_{\theta}=\mathbf{\Lambda}(\theta)$ and denote the count of non-zero elements in a matrix $\mathbf{M}$ as $[\mathbf{M}]$. For example, for a simple death process starting from $N$ members, we have $N$ non-zero elements.
Such systems in discrete space evolving in continuous time with rate matrices are termed MJPs. An MJP’s dynamics in a time interval $[0,T]$, usually thought of as an experimental time-course, are described completely by the states the system occupies over time, $\tilde{s}_{0},\tilde{s}_{1},\ldots,\tilde{s}_{l}$ with each of these occupied states $\tilde{s}_{l^{\prime}}$ being one of the state space elements, $\tilde{s}_{l^{\prime}}\in\sigma_{1:N}$, and the time at which each transition occurs, $t_{1},t_{2},\ldots,t_{l}$. Together, these describe a trajectory, an equivalent description of the MJP’s dynamics as function $\mathcal{T}(\cdot)$ mapping times from the interval $[0,T]$ onto the state space
@@ 2.1.1 Forward Modeling
Now, sampling trajectories from (10) yields MJP trajectories mediated by $\mathbf{A}_{\theta}$, described in Sec 2.1. As proven in Ref. [27], the jump times themselves may be uniformly sampled across $[0,t]$, resulting in using the term “uniformization” to describe the use of (10) in the simulation of MJPs [27, 28, 3]. Unlike the Gillespie simulation, $k$ also includes self-transitions for uniformization to remain equivalent to its corresponding MJP under uniform jump time statistics. We therefore follow the uniformization literature and say $k$ counts ‘virtual jumps’ (i.e. it is distinct from the number of transitions $l$ in the previous section), to indicate that it includes both real and self-transitions. Although uniformization is unpopular relative to Gillespie’s more efficient stochastic simulation algorithm [37], we describe it here since it is the standard in proposing trajectories conditioned on data subsequently accepted or rejected within Monte Carlo.
@@ 2.1.2 Inverse Modeling
With the details of forward modeling established, we turn to the task, the so-called inverse problem, requiring the probability of states $\rho_{n}$, given $\theta$, usually by matrix exponentiation. In the inverse problem, information on $\theta$ is to be decoded from observations. Notably, Methods #1 and #2 provide stochastic samples of trajectories and jumps, respectively, contributing to the matrix exponential while Methods #3 – #5, in principle, can be used to compute the vector times matrix exponential directly as in (2). Later, Sec. 3.2, we will embed all methods within an inverse modeling scheme as a means of assessing the efficiency of each method. In order to do so, we immediately consider multiple back to back intervals, $[0,T]$. We collect the end times of each interval as the set $\bar{T}=\{T^{1},T^{2},\ldots,T^{I}\}$. The state of a system could be observed at these times leading to a sequence of state observations $\bar{s}=\{s^{1},s^{2},\ldots,s^{I}\}$. Alternately, we could also consider a different scheme involving a collection of systems at times $\bar{T}=\{T^{1},T^{2},\ldots,T^{I}\}$ where each system’s state is observed at each of the times also yielding a sequence of state observations $\bar{s}=\{s^{1},s^{2},\ldots,s^{I}\}$. The latter falls within the realm of what is sometimes called snapshot data[9]. The methods we discuss here apply to both types of observations though, for ease of computation, our benchmarking of Methods #1 – #5 is performed on snapshot data.
Following notation established in (3), from the state observations we write $p(s^{i}|\theta)=p(\mathcal{T}^{i}(T^{i})=s^{i}|\theta)$, often termed the likelihood. As defined in (4), we have $p(s^{i}|\theta)=\rho(T^{i})_{s^{i}}=(\bar{\rho}(0)e^{\mathbf{A}_{\theta}T^{i}})_{s^{i}}$, equivalent to what is obtained by the vector times matrix exponential in (2).
In the following section we will explain different ways to avoid the full matrix exponential calculation, but first, we establish the notation necessary to explain Methods #1 – #3. Learning the kinetic parameters $\theta$ from the data means obtaining $p(\theta|\bar{s})$ related to the likelihood through Bayes’ theorem
@@
For Methods #1 and #2, we do not directly compute $p(\theta|\bar{s})$ but rather introduce latent variables $\mathcal{T}^{i}$, the trajectories, and $k^{i}$, the number of virtual jumps. That is, we re-write the posterior we desire as arising from the marginalization over trajectories and virtual jumps, respectively,
@@ 2.2 Avoiding the matrix exponential in Methods #1 – #5
We obtain concise time and memory scaling between different methods under two assumptions inspired by relevant physical examples: a) the number of nonzero elements in $\mathbf{A}_{\theta}$ scales linearly with $N$ (conveniently guaranteeing sparse $\mathbf{A}_{\theta}$) and b) the exponent of any power of $N$ appearing explicitly in $\mathbf{A}_{\theta}$ is less than or equal to $1$ as larger rates can introduce more jumps in the trajectories introduced in Methods #1 – #3 and reduce the time step of integration of Methods #4 and #5. To be clear, both conditions are used to describe how the time and memory required to calculate likelihood scales with the number of particles. They are not, however, required for the methods to work.
Regardless of which method we are discussing, when dealing with chemical reactions, the number of nonzero elements scales at worst linearly with the state space size, $O([\mathbf{A}_{\theta}])=O(N)$, as in condition a). It is also usually the case that when molecular species do not interact with themselves, the largest element of the rate matrix scales linearly with the state space — $O(\max\limits_{n}{|a_{nn}|})=O(N)$, as in condition b). As we will see, all methods have some dependence on the largest value in the rate matrix, $\max\limits_{n}{|a_{nn}|}$. As this quantity represents the fastest time-scale at which dynamics occur, all methods must take such scale into account in order to provide accurate results.
Table 1: Table summarizing the methods discussed, their number of iterations (first column) and time cost per iteration (second column), with $\kappa$ being the number of vectors taken into the Krylov subspace. The third column (time scaling) is built by multiplying the first and second columns, assuming conditions a) and b) outlined in the second paragraph of Sec. 2.2. Although J-MJP appears competitive with other methods, we will see in Sec. 3.2.2 that the latent variables it requires lead to poor MCMC mixing. The variable $k^{\ast}$ related to the R-MJP method (#3) is properly defined in Sec. 2.2.3, here it is important to know it scales proportionally to $\max\limits_{n}{|a_{nn}|}$.